Pka of the acid is equal to pH you get at the half equivalence point (4.9). Find the Ka of the acid with the following equation....
pka=4.9

-log Ka= 4.9
log Ka= -4.9

Ka=?
Answer #1

log ka = -4.9

Ka = antilog ( -4.9 )

= 1.2 * 10 -5 ( answer )

For antilog calculation, [ use next integer term ]

-4.9 = -5 ( it is used as power of 10 always ) + 0.1 (think it as log x=0.1

Then antilog value will be X * 10 ^ power